### v1

@@ Abstract
Vision Transformers (ViTs) have gained prominence as a preferred choice for a wide range of computer vision tasks due to their exceptional performance. However, their widespread adoption has raised concerns about security in the face of malicious attacks. Most existing methods rely on empirical adjustments during the training process, lacking a clear theoretical foundation. In this study, we address this gap by introducing SpecFormer, specifically designed to enhance ViTs’ resilience against adversarial attacks, with support from carefully derived theoretical guarantees. We establish local Lipschitz bounds for the self-attention layer and introduce a novel approach, Maximum Singular Value Penalization (MSVP), to attain precise control over these bounds. We seamlessly integrating MSVP into ViTs’ attention layers, using the power iteration method for enhanced computational efficiency. The modified model, SpecFormer, effectively reduces the spectral norms of attention weight matrices, thereby enhancing network local Lipschitzness. This, in turn, leads to improved training efficiency and robustness. Extensive experiments on CIFAR and ImageNet datasets confirm SpecFormer’s superior performance in defending against adversarial attacks.
@@ 1 Introduction
Vision Transformer (ViT) (Dosovitskiy et al., 2020) has gained increasing popularity in computer vision. Owing to its superior performance, ViT has been widely applied to image classification (Xu et al., 2022), object detection (Carion et al., 2020), semantic segmentation (Strudel et al., 2021), and video understanding (Arnab et al., 2021). More recently, the popular vision foundation models (Radford et al., 2021; Dehghani et al., 2023) also adopt ViT as the basic module. Unlike CNNs (Simonyan and Zisserman, 2014), a classical vision backbone, ViT initially divides images into non-overlapping patches and leverages the self-attention mechanism (Vaswani et al., 2017) for feature extraction.
@@
Despite its popularity, security concerns related to ViT have recently surfaced as a critical issue. Studies have demonstrated that ViT is vulnerable to malicious attacks (Fu et al., 2022; Lovisotto et al., 2022), which compromises its performance and system security. Adversarial examples, which are created by adding trainable perturbations to the original inputs to produce incorrect outputs, are one of the major threats in machine learning security. Different attacks, such as FGSM (Goodfellow et al., 2015), PGD (Madry et al., 2018), and CW attack (Carlini and Wagner, 2017), have significantly impeded neural networks including both CNNs (Simonyan and Zisserman, 2014; Szegedy et al., 2015; He et al., 2016) and ViT (Dosovitskiy et al., 2020; Touvron et al., 2021a; d’Ascoli et al., 2021).
@@
This work aims at improving the adversarial robustness of ViT. Prior arts (Bhojanapalli et al., 2021; Naseer et al., 2021; Bai et al., 2021b) have empirically investigated the intrinsic robustness of ViT compared with CNNs. Bhojanapalli et al. (2021) found out that when pre-trained with a sufficient amount of data, ViTs are at least as robust as the CNNs (Simonyan and Zisserman, 2014) on a broad range of perturbations. Bai et al. (2021b) stated that CNNs can be as robust as ViT against adversarial attacks if they properly adopt Transformers’ training recipes. Furthermore, Shao et al. (2021); Paul and Chen (2022) used frequency filters to discover that the success of ViT’ robustness lies in its lower sensitivity to high-frequency perturbations. But later works (Fu et al., 2022; Lovisotto et al., 2022; Wang et al., 2022b) suggest that adding an attention-aware loss to ViTs can manipulate its output, resulting in lower robustness than CNNs.
@@
Other than the robustness analysis of ViT, there are some empirical and theoretical efforts in designing algorithms to enhance its robustness. Empirically, Debenedetti et al. (2022) proposed a modified adversarial training recipe of more robust ViT by omitting the heavy data augmentations used in standard training. Mo et al. (2022) found that masking gradients from attention blocks or masking perturbations on some patches during adversarial training can greatly improve the robustness of ViT. While prior arts have mainly focused on improvements from an empirical perspective, they do not pay attention to the underlying theoretical principles governing self-attention and model robustness.
@@
Some recent research (Zhou et al., 2022; Wang et al., 2022a; Qi et al., 2023) have investigated the theoretical properties of the self-attention mechanisms in ViTs. However, they focused on studying the stability during training (Wang et al., 2022a; Takase et al., 2022) without establishing an explicit link to robustness. While some scattered understandings about the robustness of attention mechanisms do exist, their applicability and reliability remain unclear. Therefore, it is imperative to conduct a comprehensive exploration on the robustness of ViT from both the theoretical and empirical perspectives. By doing so, we can possess a more thorough understanding of what influences the robustness of ViT, which is essential for the continued development of powerful and reliable deep learning models, especially large foundation models.
@@
In this work, we propose SpecFormer, a simple yet effective approach to enhance the adversarial robustness of ViT. Concretely, we first provide a rigorous theoretical analysis of model robustness from the perspective of Lipschitz continuity (Murdock, 1999). Our analysis shows that we can easily control the Lipschitz continuity of self-attention by adding additional penalization. SpecFormer seamlessly integrate our proposed Maximum Singular Value Penaliztaion (MSVP) algorithm into each attention layer to help improve model stability. We further adopt the power iteration algorithm (Burden et al., 2015) to accelerate optimization. Our approach is evaluated on four public datasets, namely, CIFAR-10/100 (Krizhevsky et al., 2009), ImageNet (Deng et al., 2009) and Imagenette (Howard, 2019) across different ViT variants. Our SpecFormer achieves superior performance in both clean and robust accuracy. Under standard training, our approach improves robust accuracy by $\mathbf{2.79}$and $\mathbf{2.89}$against FGSM (Goodfellow et al., 2015) and PGD (Madry et al., 2018) attacks, respectively; under adversarial training, SpecFormer outperforms the best counterparts by $\mathbf{3.35}$and $\mathbf{2.58}$on CW (Carlini and Wagner, 2017) and PGD (Madry et al., 2018) attacks. The clean accuracy is also improved by $\mathbf{1.67}\%,\text{ and }\mathbf{3}\%$ on average in both settings.
Our main contributions can be summarized as follows:
  • We provide a comprehensive theoretical analysis of the local Lipschitz continuity of Vision Transformer robustness and compare it with existing bounds in the literature.
  • We propose SpecFormer, a robust Transformer architecture that applies Maximum Spectral Value Penalization (MSVP) to each attention layer to enhance adversarial robustness.
  • Extensive experiments demonstrate the superiority of our approach in both clean and robust accuracy against different adversarial attacks.
@@ 2.1 Adversarial Robustness
The early work of adversarial robustness (Szegedy et al., 2013) discovered that although deep networks are highly expressive, their learned input-output mappings are fairly discontinuous to some extent. Therefore, people can apply an imperceptible perturbation that maximizes the network’s prediction error to cause misclassification of images. Since then, a vast amount of research (Goodfellow et al., 2015; Nguyen et al., 2015; Papernot et al., 2016a) has been done in adversarial attack (Carlini and Wagner, 2017; Chen et al., 2018), defense (Papernot et al., 2016b; Xie et al., 2018), robustness (Zheng et al., 2016), and theoretical understanding (Pang et al., 2022; Zhang et al., 2019a).
The problem of overall adversarial robustness can be formulated as a min-max optimization, $\min_{\bm{\theta}}\max_{\bm{\delta}}\mathcal{L}(f(\mathbf{x}+\bm{\delta}),y)$. The solution to the inner maximization problem w.r.t the input perturbations $\bm{\delta}$ corresponds to generating adversarial samples (Goodfellow et al., 2015; Madry et al., 2018), while the solution to the outer minimization problem w.r.t the model parameters $\bm{\theta}$ corresponds to adversarial training (Tramèr et al., 2018; Bai et al., 2021a), which is an irreplaceable method for improving adversarial robustness and plays a crucial role in defending against adversarial attacks.
Existing attack approaches focus on deriving more and more challenging perturbations $\bm{\delta}$ to explore the limits of neural network adversarial robustness. For instance, Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2015) directly take the sign of the derivative of the training loss w.r.t the input perturbation as $\bm{\delta}\leftarrow\bm{\delta}+\alpha\operatorname{sign}\left(\nabla_{\bm{\delta}}\mathcal{L}(f_{\theta}(\mathbf{x}+\bm{\delta}),y)\right)$. Projected Gradient Descent (PGD) method (Madry et al., 2018) updates the perturbations by taking the sign of the loss function derivative w.r.t the perturbation, projecting the updated perturbations onto the admissible space, and repeating multiple times to generate more powerful attacks: $\bm{\delta}\leftarrow\Pi_{\epsilon}\left(\bm{\delta}+\alpha\cdot\operatorname{sign}\left(\nabla_{\bm{\delta}}\mathcal{L}\left(f_{\bm{\theta}}(\mathbf{x}+\bm{\delta}),y\right)\right)\right)$.
@@ 2.2 Robustness of Vision Transformer
Regarding the robustness of Vision Transformers, the first question people are curious about is how they compare to the classic CNN structure in terms of robustness. To address this question, numerous papers (Szegedy et al., 2013; Shao et al., 2021; Bhojanapalli et al., 2021; Bai et al., 2021a) have conducted thorough evaluations of both visual model structures, analyzed possible causes for differences in robustness, and proposed solutions to mitigate the observed gaps. Building upon these understandings, Zhou et al. (2022) developed fully attentional networks (FANs) to enhance the robustness of self-attention mechanisms. They evaluate the efficacy of these models in terms of corruption robustness for semantic segmentation and object detection, achieving state-of-the-art results. Additionally, Debenedetti et al. (2022) compared the DeiT (Touvron et al., 2021a), CaiT (Touvron et al., 2021b), and XCiT (Ali et al., 2021) ViT variants in adversarial training and discovered that XCiT was the most effective. This discovery sheds light on the idea that Cross-Covariance Attention could be another viable option for improving the adversarial robustness of ViTs.
In terms of the theoretical understanding of the relationship between Transformers and robustness, Zhou et al. (2022) attribute the emergence of robustness in Vision Transformers to the connection of self-attention mechanisms with information bottleneck theory. This suggests that the stacking of attention layers can be broadly regarded as an iterative repeat of solving an information-theoretic optimization problem, which promotes grouping and noise filtering. Dasoulas et al. (2021) analyzed that the norm of the derivative of attention models is directly related to the uniformity of the softmax probabilities. If all attention heads have uniform probabilities, the norm will reach its minimum; if the whole mass of the probabilities is on one element, the norm will reach its maximum. Considering a linear approximation of the attention mechanism, the smaller the derivative norm, the tighter the changes when perturbations are introduced, thus enhancing the robustness of the Transformer. These findings have been supported by another paper (Lovisotto et al., 2022) that proposes an attention-aware loss to deceive the predictions of Vision Transformers. Their attack strategy exactly misguides the attention of all queries towards a single key token under the control of an adversarial patch, corresponding to the maximum derivative norm case described above.
@@ 3 Preliminaries
Denote a clean dataset as $\mathcal{D}=\{\mathbf{x}_{i},y_{i}\}_{i=1}^{N}$, where $\mathbf{x}$ and $y$ are the input and ground-truth label respectively, and denote the loss function as $\mathcal{L}$. We aim to investigate the robustness of the Vision Transformer (ViT) under two different training paradigms: standard supervised training and adversarial training (AT) (Madry et al., 2018). In standard training, we aim to learn a classification function $f_{\bm{\theta}}$ parameterized by $\bm{\theta}$, where the optimal parameter is $\bm{\theta}^{*}=\operatorname{\arg\min}_{\bm{\theta}}\mathds{E}_{(\mathbf{x},y)\in\mathcal{D}}\leavevmode\nobreak\ \mathcal{L}(f_{\bm{\theta}}(\mathbf{x}),y)$.
While standard training is shown to be less resilient to adversarial attacks, AT is a major and effective paradigm for adversarial robustness. AT aims to improve a model’s ability to resist malicious attacks by solving a min-max optimization problem and generating adversarial examples by adding strategic perturbations $\bm{\delta}$ within a given budget $\delta_{0}$, which can fool the model predictions, i.e., $f_{\bm{\theta}}(\mathbf{x}+\bm{\delta})\neq y$. The model parameters are updated to reduce the classification error on the perturbed examples, resulting in enhanced robustness against adversarial attacks:
@@
where $\delta_{0}$ bounds the magnitude of $\bm{\delta}$ to prevent unintended semantic changes caused by perturbation.
@@ 4 Theoretical Analysis
Bounding the global Lipschitz constant of a neural network is a commonly used method to provide robustness guarantees (Cisse et al., 2017; Leino et al., 2021). However, the global Lipschitz bound can be loose because it needs to hold for all points in the input domain, including inputs that are far apart. This can greatly reduce clean accuracy in empirical comparisons (Huster et al., 2019; Madry et al., 2018). On the other hand, a local Lipschitz constant bounds the norm of output perturbations only for inputs within a small region, typically selected as a neighborhood around each data point. This aligns perfectly with the scenario of adversarial robustness, as discussed in Section 2.1, where perturbations attempt to affect the model’s output within a budget constraint. Local Lipschitz bounds are superior because they produce tighter bounds by considering the geometry in a local region, often leading to much better robustness (Hein and Andriushchenko, 2017; Zhang et al., 2019b).
The core of our study is to bridge the gap between local Lipschitz continuity and adversarial robustness in ViTs. Primary distinctions of ViT lie in the LayerNorm and self-attention layers. The seminal work (Kim et al., 2021) proved the dot-product is not globally Lipschitz continuous and the LayerNorm is Lipschitz continuous (see Kim et al. (2021), Appendix N). Therefore, we only need to modify that concept and prove that the dot-product self-attention is local Lipschitz continuous. By utilizing the local Lipschitz continuity, the adversarial robustness can be strengthened.
Definition 4.1 (Local Lipschitz Continuity).
@@
Definition 4.2 (Local Lipschitz Constant).
@@
Theorem 4.1 (Calculation of Local Lipschitz Constant (Federer, 1969) ).
@@
The global Lipschitz constant, which takes into account the supremum over ${\mathcal{X}}={\mathbb{R}}^{d}$, must ensure Eq. (2) even for distant ${\mathbf{x}}$ abd ${\mathbf{x}}_{0}$, which can render it imprecise and lacking significance when examining the local behavior of a network around a single input. We concentrate on $2$-local Lipschitz constants, where ${\mathcal{X}}=B_{2}({\mathbf{x}}_{0},\delta_{0})\coloneqq\{{\mathbf{x}}:\|{\mathbf{x}}-{\mathbf{x}}_{0}\|_{2}\leq\delta_{0}\}$ represents a small $\ell_{2}$-ball with a radius of $\delta_{0}$ centered around ${\mathbf{x}}_{0}$. The choice of the $2$-norm for our investigation is motivated by two key reasons. First, the $2$-norm is the most commonly used norm in Euclidean space. Second, and delving further into the rationale, as revealed in (Yoshida and Miyato, 2017), the sensitivity of a model to input perturbations is intricately connected to the $2$-norm of the weight matrices, which in turn is closely linked to the principal direction of variation and the maximum scale change.
Proposition 4.2 (Connection of Model Sensitivity and Maximum Singular Value (Yoshida and Miyato, 2017)).
@@
$h_{1}(\mathbf{X})=\mathbf{X}\mathbf{W}^{Q},\leavevmode\nobreak\ h_{2}(\mathbf{X})=\mathbf{X}\mathbf{W}^{K},\leavevmode\nobreak\ h_{3}(\mathbf{X})=\mathbf{X}\mathbf{W}^{V}.$ (8)
After reinterpreting the self-attention mechanism as a product of three linear mappings, we are inspired by Prop. 4.2 to readily discern that we can independently control the maximum singular values of these three weight matrices to imbue the model with adversarial robustness.
Theorem 4.3.
@@
${\operatorname{Lip}_{local}(\mathrm{Att},\mathbf{X}_{0})}\leq N(N+1)\left(\|X_{0}\|_{F}+\delta_{0}\right)^{2}\Big{[}\left\|\mathbf{W}^{V}\right\|_{2}\left\|\mathbf{W}^{Q}\right\|_{2}\left\|\mathbf{W}^{K,\top}\right\|_{2}+\left\|\mathbf{W}^{V}\right\|_{2}\Big{]},$ (9)
@@
${\operatorname{Lip}_{local}(\mathrm{Att},\mathbf{X}_{0})}\leq N(N+1)\left(B+\delta_{0}\right)^{2}\Big{[}\left\|\mathbf{W}^{V}\right\|_{2}\left\|\mathbf{W}^{Q}\right\|_{2}\left\|\mathbf{W}^{K,\top}\right\|_{2}+\left\|\mathbf{W}^{V}\right\|_{2}\Big{]}.$ (10)

### v2

@@ Abstract
Vision Transformers (ViTs) are increasingly used in computer vision due to their high performance, but their vulnerability to adversarial attacks is a concern. Existing methods lack a solid theoretical basis, focusing mainly on empirical training adjustments. This study introduces SpecFormer, tailored to fortify ViTs against adversarial attacks, with theoretical underpinnings. We establish local Lipschitz bounds for the self-attention layer and propose the Maximum Singular Value Penalization (MSVP) to precisely manage these bounds By incorporating MSVP into ViTs’ attention layers, we enhance the model’s robustness without compromising training efficiency. SpecFormer, the resulting model, outperforms other state-of-the-art models in defending against adversarial attacks, as proven by experiments on CIFAR and ImageNet datasets. Code is released at https://github.com/microsoft/robustlearn.
Keywords:
Vision Transformer Adversarial Robustness Lipschitz Continuity
@@ 1 Introduction
Vision Transformer (ViT) [17] has gained increasing popularity in computer vision. Owing to its superior performance, ViT has been widely applied to image classification [58], object detection [8], semantic segmentation [46], and video understanding [2]. More recently, the popular vision foundation models [43, 15] also adopt ViT as the basic module. Unlike CNNs [45], a classical vision backbone, ViT initially divides images into non-overlapping patches and leverages the self-attention mechanism [53] for feature extraction.
@@
Despite its popularity, security concerns related to ViT have recently surfaced as a critical issue. Studies have demonstrated that ViT is vulnerable to malicious attacks [20, 30], which compromises its performance and system security. Adversarial examples, which are created by adding trainable perturbations to the original inputs to produce incorrect outputs, are one of the major threats in machine learning security. Different attacks, such as FGSM [21], PGD [32], and CW attack [9], have significantly impeded neural networks including both CNNs [45, 47, 22] and ViT [17, 50, 18].
@@
This work aims at improving the adversarial robustness of ViT. Prior arts [6, 36, 4] have empirically investigated the intrinsic robustness of ViT compared with CNNs. [6] found out that when pre-trained with a sufficient amount of data, ViTs are at least as robust as the CNNs [45] on a broad range of perturbations. [4] stated that CNNs can be as robust as ViT against adversarial attacks if they properly adopt Transformers’ training recipes. Furthermore, [44, 41] used frequency filters to discover that the success of ViT’ robustness lies in its lower sensitivity to high-frequency perturbations. But later works [20, 30, 56] suggest that adding an attention-aware loss to ViTs can manipulate its output, resulting in lower robustness than CNNs.
@@
Other than the robustness analysis of ViT, there are some empirical and theoretical efforts in designing algorithms to enhance its robustness. Empirically, [14] proposed a modified adversarial training recipe of more robust ViT by omitting the heavy data augmentations used in standard training. [34] found that masking gradients from attention blocks or masking perturbations on some patches during adversarial training can greatly improve the robustness of ViT. While prior arts have mainly focused on improvements from an empirical perspective, they do not pay attention to the underlying theoretical principles governing self-attention and model robustness.
@@
Some recent research [63, 55, 42] have investigated the theoretical properties of the self-attention mechanisms in ViTs. However, they focused on studying the stability during training [55, 49] without establishing an explicit link to robustness. While some scattered understandings about the robustness of attention mechanisms do exist, their applicability and reliability remain unclear. Therefore, it is imperative to conduct a comprehensive exploration on the robustness of ViT from both the theoretical and empirical perspectives. By doing so, we can possess a more thorough understanding of what influences the robustness of ViT, which is essential for the continued development of powerful and reliable deep learning models, especially large foundation models.
@@
In this work, we propose SpecFormer, a simple yet effective approach to enhance the adversarial robustness of ViT. Concretely, we first provide a rigorous theoretical analysis of model robustness from the perspective of Lipschitz continuity [35]. Our analysis shows that we can easily control the Lipschitz continuity of self-attention by adding additional penalization. SpecFormer seamlessly integrate our proposed Maximum Singular Value Penaliztaion (MSVP) algorithm into each attention layer to help improve model stability. We further adopt the power iteration algorithm [7] to accelerate optimization. Our approach is evaluated on four public datasets, namely, CIFAR-10/100 [27], ImageNet [16] and Imagenette [24] across different ViT variants. Our SpecFormer achieves superior performance in both clean and robust accuracy. Under standard training, our approach improves robust accuracy by $\mathbf{8.96}$%, $\mathbf{3.49}$%, and $\mathbf{2.36}$against FGSM [21], PGD [32], and AutoAttack [12], respectively. Under adversarial training, SpecFormer outperforms the best counterparts by $\mathbf{1.69}$%, $\mathbf{1.26}$%, and $\mathbf{0.50}$on CW [9], PGD [32], and AutoAttack [12]. The clean accuracy is also improved by $\mathbf{2.20}\%,\text{ and }\mathbf{3.06}\%$ on average in both settings.
@@ 2.1 Adversarial Robustness
The early work of adversarial robustness [48] discovered that although deep networks are highly expressive, their learned input-output mappings are fairly discontinuous. Therefore, people can apply an imperceptible perturbation that maximizes the network’s prediction error to cause misclassification. Since then, a vast amount of research [21, 37, 39] has been done in adversarial attack [9, 10], defense [40, 57], robustness [62], and theoretical understanding [38, 60].
The problem of overall adversarial robustness can be formulated as a min-max optimization, $\min_{\bm{\theta}}\max_{\bm{\delta}}\mathcal{L}(f(\mathbf{x}+\bm{\delta}),y)$. The solution to the inner maximization problem w.r.t the input perturbations $\bm{\delta}$ corresponds to generating adversarial samples [21, 32], while the solution to the outer minimization problem w.r.t the model parameters $\bm{\theta}$ corresponds to adversarial training [52, 3], which is an irreplaceable method for improving adversarial robustness and plays a crucial role in defending against adversarial attacks.
Existing attack approaches focus on deriving more and more challenging perturbations $\bm{\delta}$ to explore the limits of neural network adversarial robustness. For instance, Fast Gradient Sign Method (FGSM) [21] directly take the sign of the derivative of the training loss w.r.t the input perturbation as $\bm{\delta}\leftarrow\bm{\delta}+\alpha\operatorname{sign}\left(\nabla_{\bm{\delta}}\mathcal{L}(f_{\theta}(\mathbf{x}+\bm{\delta}),y)\right)$. Projected Gradient Descent (PGD) method [32] updates the perturbations by taking the sign of the loss function derivative w.r.t the perturbation, projecting the updated perturbations onto the admissible space, and repeating multiple times to generate more powerful attacks:
$\bm{\delta}\leftarrow\Pi_{\epsilon}\left(\bm{\delta}+\alpha\cdot\operatorname{sign}\left(\nabla_{\bm{\delta}}\mathcal{L}\left(f_{\bm{\theta}}(\mathbf{x}+\bm{\delta}),y\right)\right)\right).$
@@ 2.2 Robustness of Vision Transformer
Regarding the robustness of Vision Transformers, the first question people are curious about is how they compare to the classic CNN structure in terms of robustness. To address this question, numerous papers [48, 44, 6, 3] have conducted thorough evaluations of both visual model structures, analyzed possible causes for differences in robustness, and proposed solutions to mitigate the observed gaps. Building upon these understandings, [63] developed fully attentional networks (FANs) to enhance the robustness of self-attention mechanisms. They evaluate the efficacy of these models in terms of corruption robustness for semantic segmentation and object detection, achieving state-of-the-art results. Additionally, [14] compared the DeiT [50], CaiT [51], and XCiT [1] ViT variants in adversarial training and discovered that XCiT was the most effective. This discovery sheds light on the idea that Cross-Covariance Attention could be another viable option for improving the adversarial robustness of ViTs.
In terms of the theoretical understanding of the relationship between Transformers and robustness, [63] attribute the emergence of robustness in Vision Transformers to the connection of self-attention mechanisms with information bottleneck theory. This suggests that the stacking of attention layers can be broadly regarded as an iterative repeat of solving an information-theoretic optimization problem, which promotes grouping and noise filtering. [13] analyzed that the norm of the derivative of attention models is directly related to the uniformity of the softmax probabilities. If all attention heads have uniform probabilities, the norm will reach its minimum; if the whole mass of the probabilities is on one element, the norm will reach its maximum. Considering a linear approximation of the attention mechanism, the smaller the derivative norm, the tighter the changes when perturbations are introduced, thus enhancing the robustness of the Transformer. These findings have been supported by another paper [30] that proposes an attention-aware loss to deceive the predictions of Vision Transformers. Their attack strategy exactly misguides the attention of all queries towards a single key token under the control of an adversarial patch, corresponding to the maximum derivative norm case described above.
@@ 3 Preliminaries
Denote a clean dataset as $\mathcal{D}=\{\mathbf{x}_{i},y_{i}\}_{i=1}^{N}$, where $\mathbf{x}$ and $y$ are the input and ground-truth label respectively, and denote the loss function as $\mathcal{L}$. We aim to investigate the robustness of the Vision Transformer (ViT) under two different training paradigms: standard supervised training and adversarial training (AT) [32]. In standard training, we aim to learn a classification function $f_{\bm{\theta}}$ parameterized by $\bm{\theta}$, where the optimal parameter is $\scriptstyle\bm{\theta}^{*}=\operatorname{\arg\min}_{\bm{\theta}}\mathds{E}_{(\mathbf{x},y)\in\mathcal{D}}~{}\mathcal{L}(f_{\bm{\theta}}(\mathbf{x}),y)$.
While standard training is shown to be less resilient to adversarial attacks, adversarial training (AT) is a major and effective paradigm for enhancing adversarial robustness. AT aims to improve a model’s ability to resist malicious attacks by solving a min-max optimization problem. For the inner optimization, it generates adversarial examples that are as powerful as possible to deceive the model predictions, i.e., $f_{\bm{\theta}}(\mathbf{x}+\bm{\delta})\neq y$. For the outer optimization, it adjusts the model parameters to minimize the classification error caused by the perturbed examples, resulting in enhanced robustness:
@@
where $\delta_{0}$ bounds the magnitude of $\bm{\delta}$ to prevent unintended semantic changes caused by the perturbation.
@@ 4 Theoretical Analysis
Bounding the global Lipschitz constant of a neural network is a commonly used method to provide robustness guarantees [11, 28]. However, the global Lipschitz bound can be loose because it needs to hold for all points in the input domain, including inputs that are far apart. This can greatly reduce clean accuracy in empirical comparisons [25, 32]. Conversely, a local Lipschitz constant bounds the norm of output perturbations for inputs within a small region, typically selected as a neighborhood around each data point. This aligns perfectly with the scenario of adversarial robustness, as discussed in Section 2.1, where perturbations attempt to affect the model’s output within a budget constraint. Local Lipschitz bounds are superior because they produce tighter bounds by considering the geometry in a local region, often leading to much better robustness [23, 61].
The core of our study is to bridge the gap between local Lipschitz continuity and adversarial robustness in ViTs. Primary distinctions of ViT lie in the LayerNorm and self-attention layers. The seminal work [26] proved the dot-product is not globally Lipschitz continuous and the LayerNorm is Lipschitz continuous [26]. Therefore, we only need to modify that concept and prove that the dot-product self-attention is local Lipschitz continuous. By utilizing the local Lipschitz continuity, the adversarial robustness can be strengthened.
Definition 1 (Local Lipschitz Continuity)
@@
Definition 2 (Local Lipschitz Constant)
@@
Theorem 4.1 (Calculation of Local Lipschitz Constant [19] )
@@
The global Lipschitz constant, which takes into account the supremum over ${\mathcal{X}}={\mathbb{R}}^{d}$, must ensure Eq. (2) even for distant ${\mathbf{x}}$ abd ${\mathbf{x}}_{0}$, which can render it imprecise and lacking significance when examining the local behavior of a network around a single input. We concentrate on $2$-local Lipschitz constants, where ${\mathcal{X}}=B_{2}({\mathbf{x}}_{0},\delta_{0})\coloneqq\{{\mathbf{x}}:\|{\mathbf{x}}-{\mathbf{x}}_{0}\|_{2}\leq\delta_{0}\}$ represents a small $\ell_{2}$-ball with a radius of $\delta_{0}$ centered around ${\mathbf{x}}_{0}$. The choice of the $2$-norm for our investigation is motivated by two key reasons. First, the $2$-norm is the most commonly used norm in Euclidean space. Second, and delving further into the rationale, as revealed in [59], the sensitivity of a model to input perturbations is intricately connected to the $2$-norm of the weight matrices, which in turn is closely linked to the principal direction of variation and the maximum scale change.
Proposition 1 (Model Sensitivity and Maximum Singular Value in Linear Models[59])
@@
$h_{1}(\mathbf{X})=\mathbf{X}\mathbf{W}^{Q},~{}h_{2}(\mathbf{X})=\mathbf{X}\mathbf{W}^{K},~{}h_{3}(\mathbf{X})=\mathbf{X}\mathbf{W}^{V}.$ (8)
After reinterpreting the self-attention mechanism as a product of three linear mappings, we are inspired by Prop. 1 to readily discern that we can independently control the maximum singular values of these three weight matrices to imbue the model with adversarial robustness.
Theorem 4.2
@@
${\operatorname{Lip}_{local}(\mathrm{Att},\mathbf{X}_{0})}\leq N(N+1)\left(\|X_{0}\|_{F}+\delta_{0}\right)^{2}\Big{[}\left\|\mathbf{W}^{V}\right\|_{2}\left\|\mathbf{W}^{Q}\right\|_{2}\left\|\mathbf{W}^{K,\top}\right\|_{2}+\left\|\mathbf{W}^{V}\right\|_{2}\Big{]},$
@@
${\operatorname{Lip}_{local}(\mathrm{Att},\mathbf{X}_{0})}\leq N(N+1)\left(B+\delta_{0}\right)^{2}\Big{[}\left\|\mathbf{W}^{V}\right\|_{2}\left\|\mathbf{W}^{Q}\right\|_{2}\left\|\mathbf{W}^{K,\top}\right\|_{2}+\left\|\mathbf{W}^{V}\right\|_{2}\Big{]}.$
@@
Comparison with Existing Bounds
The existing literature on introducing Lipschitz continuity into Transformer models comprises three notable articles[26, 13, 42], where we abbreviate their proposed modified Transformers as L2Former[26], LNFormer[13], and LipsFormer[42]. We compare these four models from three perspectives: whether they introduce Lipschitz continuity in Transformers, whether they address the robustness problem, and whether the analysis they provide on the attention mechanism is simplified or not. A detailed summary is presented in Table 1. We can see that our analysis is the most comprehensive. More detailed comparisons can be found in Appendix D.
Table 1: Comprehensive Comparison of the Baseline Models.
Model Proposed Mechanism Lipschitz Continuity Robustness Not Simplified
Transformer [17] $\operatorname{softmax}\left(\frac{\mathbf{X}\mathbf{W}^{Q}\left(\mathbf{X}\mathbf{W}^{K}\right)^{\top}}{\sqrt{D}}\right)\mathbf{X}\mathbf{W}^{V}$-
L2Former [26] $\exp\left(-\frac{\left\|\mathbf{x}_{i}^{\top}W^{Q}-\mathbf{x}_{j}^{\top}W^{K}\right\|_{2}^{2}}{\sqrt{D/H}}\right)$
LNFormer [13] $\frac{Q^{\top}K}{\max\{uv,uw,vw\}}$
LipsFormer [42] $\mathbf{q}_{i}=\frac{\left(\mathbf{x}_{i}^{\top}\mathbf{W}^{Q}\right)^{\top}}{\sqrt{\left\|\mathbf{x}_{i}^{\top}\mathbf{W}^{Q}\right\|^{2}+\epsilon}}$